OpenStudy (anonymous):

if one side of a right triangle is 3 inches and the hypotenus is 4 inches, how long is the other side of the triangle? show work

OpenStudy (anonymous):

h^2 = a^2 + b^2

OpenStudy (anonymous):

A^2+B^2=C^2 Hypotenus=C=4 Side(A) =A=3 Side(B) =B=? (4^2)=(3^2)+(B^2) 16=9+B^2 16-9=B^2 7=B^2 Take the squareroot(sqrt) B=sqrt(7)